OpenStudy (anonymous):

Simplify f(x) (3x^(2)-4x+1)/(3x-1) . Using complete sentences, explain why f(1) = 0, f(0) = -1, and f(-1) = -2, yet f( ) is undefined. Make sure to show your work. (

OpenStudy (anonymous):

Factorize f(x) then U'll get \[3x ^{2}-4x+1\]= (3x-1)(x-1) f(x)= (3x-1)(x-1)/(3x-1) = x-1 Now for f(1) put x as 1 So U shoud get f(x)= x-1 f(1)= 1-1 =0 Similarly for the rest f(0) = 0-1=-1 f(-1) = -1-1=-2